Perform the operation and simplify, expressing answer with a positive exonent in lowest terms: (x^2/3)(x^1/5).
Answer by edjones(8007)

(x^2/3)(x^1/5)
Add exponents when multiplying.
2/3 + 1/5=10/15 + 3/15=13/15
x^(2/3)*x^(1/5)=x^(13/15)
Ed